Belmont Corp. is considering the purchase of a new piece of equipment. The cost savings from the equipment would result in an annual increase in net income after tax of $200,000. The equipment will have an initial cost of $1,000,000 and have an 8-year life. If there is no salvage value of the equipment, what is the payback period?

8 years

5 years

1.6 years

3.08 years
Answer #1

depreciation expense = 1000000/8 = 125000

Net cash flow = 200000+125000 = 325000

Payback period = Initial investment /Annual Cash flow = 1000000/325000 = 3.08 years

So answer is d) 3.08 years